# Cahn-Hilliard model with Schlögl Reactions: interplay of equilibrium and non-equilibrium phase transitions. II. Memory effects

### v1

@@ Abstract
The present work is the continuation of our paper [1]. It is devoted to the modeling of the interplay of equilibrium and non-equilibrium phase transitions. The modeling of equilibrium phase transition is based on the modified Cahn-Hilliard equation. The non-equilibrium phase transition is modeled by the Second Schlögl reaction system. We consider the advancing front, which combines these both transitions. Different from [1], we consider here the memory effects, i.e. the effects of non-Fickian diffusion. The traveling wave solution is obtained, its dependence on the model parameters is studied in detail. The relative importance of memory effects for different process regimes is estimated.
@@
phase transition, nonequilibrium phase transition, Cahn-Hilliard equation, Schlögl reactions, memory effect, non-Fickian diffusion, traveling wave
pacs:
64.60.A–, 64.60.De, 82.40.–g
I Introduction
@@
The present work is the continuation of our paper [1]. It was devoted to the modeling of the interplay of equilibrium and non-equilibrium phase transitions. We considered the advancing fronts which ‘‘combine’’, in some sense, these both transitions. The equilibrium phase transition was modeled on the basis of convective-viscous Cahn-Hilliard equation [2, 3]. The non-equilibrium phase transition was modeled by canonical Schlögl chemical reaction systems [4]. The insight into the history and existing modifications of the Cahn-Hilliard equation were given in [1] and will not be repeated here. We also refer to pioneering papers and excellent reviews [5, 6, 7, 8]. Below, in the introduction, we give first only brief explanation of some basic assumptions and description of the models, introduced in [1], after what we outline the specificities of the present work.
@@
The basic underlying idea of Cahn-Hilliard model is that for inhomogeneous system, e.g. system undergoing a phase transition, the thermodynamic potential (e.g. free energy) should depend not only on the order parameter $X$, but on its gradient as well. The idea of such dependence was introduced already by Van der Waals [9] in his theory of capillarity. For the inhomogeneous system the local chemical potential $\mu$ is defined as variational derivative of the thermodynamic potential functional. If the thermodynamic potential is the simplest symmetric – quadratic – function of gradient this leads to the local chemical potential which depends on Laplacian, or for the one-dimensional case – on the second order derivative of the order parameter, in our case of concentration $X$ of a component (herein our notations differ from notations in [1]):
$\mu=-\bar{\varepsilon}^{2}\Delta X+f\left(X\right).$ (1.1)
@@
$\bar{\varepsilon}$ is usually presumed to be proportional to the capillarity length, and $f\left(X\right)=\frac{d\Phi\left(X\right)}{dX}$, where $\Phi\left(X\right)$ is homogeneous part of the thermodynamic potential. In the present communication we will take $f\left(X\right)$ in the form of the cubic polynomial (corresponding to the fourth-order polynomial for the homogeneous part of thermodynamic potential):
@@
The roots of the right-hand side of (1.2) are the values of $X$ corresponding to two stable minima and one unstable maximum of the potential polynomial. In the present paper $X$ is presumed to be non-negative, so even for a symmetric potential $\bar{\delta}\neq 0$; furthermore, the asymmetric potential naturally appears in some modifications of the Cahn-Hilliard equation [8]. In this phenomenological model we consider always the isothermal situation, so we do not show the temperature dependence of the coefficients in (1.2) explicitly. But if we want to model the approach to critical state, for such a model the approach to critical temperature will be manifested by merging the stationary states together [1], i.e. by two non-zero roots of the right-hand side of (1.2) approaching the third zero root.
@@
yields, with expression (1.1) for the chemical potential, instead of usual second order diffusion equation a fourth-order PDE for the order parameter $X$:
@@
In presence of the external field the convective term is added to equation(1.5), and to account for dissipation the viscous term is added to chemical potential $\mu$ [11, 12, 13, 14, 2, 3]:
@@
Here $\bar{\alpha}$ is proportional to the field and $\bar{\eta}$ is viscosity; we consider the one-dimensional problem herein.
@@
Among numerous modifications of the Cahn-Hilliard equation we are interested here in the models, where the nonlinear sink/source terms, e.g. due to a chemical reaction, are inserted into the right-hand sides of (1.4) (and, correspondingly of (1.5) as well):
$\frac{\partial X}{\partial t^{\prime}}=-\nabla J+R\left(X\right).$ (1.7)
Such study was pioneered by Huberman [15]. He introduced Cahn-Hilliard equation with additional kinetic terms corresponding to the reversible first-order autocatalytic chemical reaction and analyzed linear stability of the stationary states. Cohen and Murray [16] considered the same equation in the biological context: they used quadratic nonlinearity to describe growth and dispersal in population model. They studied the stability and identified bifurcations to spatial structures. Similar equation (with additional nonlinear term) was used in [17] to study segregation dynamics of binary mixtures coupled with chemical reaction. The same equation as in [15, 16] was used to describe phase transitions in chemisorbed layer [18] and to model the system of cells that move, proliferate and interact via adhesion [19]. Also, for the latter model several rigorous mathematical results on existence and asymptotics of solutions were obtained [20, 21]. General observation is that the presence of chemical reaction can visibly influence the equilibrium phase transition, e.g. freeze the spinodal decomposition or coarsening, stabilizing some stationary inhomogeneous state. The general and even coordinate-dependent nonlinearity was recently considered by Miranville [22].
@@
On the other hand, the canonical models for non-equilibrium phase transitions in chemical reaction systems were introduced by Schlögl [4]; here the different ‘‘phases’’ correspond to different stationary states of the system. Schlögl considered two reaction systems: the so-called ‘‘First Schlögl Reaction’’
$A+X\rightleftarrows 2X,$ (1.8)
$B+X\rightleftarrows C$ (1.9)
and the ‘‘Second Schlögl Reaction’’
$A+2X\rightleftarrows 3X,$ (1.10)
$B+X\rightleftarrows C.$ (1.11)
Below we denote the concentrations with the same letters, as species; the concentrations of species $A$,$B$ and $C$ (which are called the ‘‘reservoir reagents’’) are presumed to be constant and only concentration of $X$ can vary with time and space. For the first Schlögl reaction in the absence of diffusion the evolution of $X$ is described by
$\frac{dX}{dt}=\,-k^{\prime}_{11}X^{2}+k_{11}AX-k_{21}BX+k^{\prime}_{21}C.$ (1.12)
Here $k_{ij}\,,\,\,k^{\prime}_{ij}\,$ are the rate constants for the forward and reverse reactions, respectively; the second lower index is ‘‘1’’ for the first Schlögl reaction, and ‘‘2’’ for the second. Correspondingly, for the second Schlögl reaction in the absence of diffusion the evolution of $X$ is described by
$\frac{dX}{dt}=\,-k^{\prime}_{12}X^{3}+k_{12}AX^{2}-k_{22}BX+k^{\prime}_{22}C.$ (1.13)
The first reaction exhibits a non-equilibrium phase transition of second order, the second one a phase transition of first order (for the details see [4]). If the system simultaneously undergoes an equilibrium phase transition accompanied by phase separation, it could be of considerable interest to study the interaction of an equilibrium and non-equilibrium phase transitions. Apparently unaware of Schlögl paper Huberman [15] and Cohen and Murray [16] in fact considered interplay of equilibrium and (second-order) non-equilibrium phase transitions.
@@
In [1] we have considered the convective-viscous Cahn-Hilliard equation (1.6), see [2, 3], complemented by source/sink terms corresponding both to first and second Schlögl reactions. We have called these modifications Cahn-Hilliard-Huberman-Cohen-Murray (CHHCM) and Cahn-Hilliard-Schlögl (CHS) equations respectively. We obtained exact traveling-wave solutions for these modifications. We gave also exact traveling-wave solution for Puri-Frish modification [17]. For convective-viscous equation with first Schlögl reaction the conditions of simultaneous equilibrium and non-equilibrium phase transitions are very restrictive and rigid. Even more, equation for the traveling wave becomes degenerate in the absence of the applied field, splitting into second and first order equations. So, the constant-velocity combined-transition-front model is not very instructive in this case. On the other hand, for the CHS the effect of the non-equilibrium transition, i.e. of the reactive system, is much stronger. The transition front may be stopped, or even reversed both by changing the stationary states of the reaction system and by the field. So, exploring the effect of the ‘‘non-Fickian’’ diffusion we will consider the convective-viscous CHS only.
@@
Indeed, in all models described above the diffusional flux was presumed to be proportional to the gradient of the chemical potential. This ‘‘modified’’ Fick’s law (the original Fick’s law presumes proportionality to the gradient of concentration) was often criticized for the infinite speed of the spread of diffusing substance. The most popular alternative is the Maxwell-Cattaneo approach, see [23] for a comprehensive discussion; here we give only the one-dimensional formulae, in accord with the spirit of the present paper. In Maxwell-Cattaneo approach the mass-conservation, or continuity equations (1.4) or (1.7) are not changed. However instead of (1.3) the following relation is proposed:
@@
[…]
So this approach is also called ‘‘diffusion with memory effects’’. Correspondingly, $\tau^{\prime}$ is considered as characteristic time of the memory. On the other hand, eliminating $J$ from (1.7) and (1.14) yields the ‘‘hyperbolic modification’’ of the diffusion equation
@@
The hyperbolic modification of the convective-viscous Cahn-Hilliard equation (1.6) with source/sink term takes the form
@@
The hyperbolic modification of the classical Cahn-Hilliard equation was proposed in [24] to model rapid spinodal decomposition in a binary alloy. However, from purely mathematical viewpoint – as a singular perturbation of the classic Cahn-Hilliard equation – it was considered even earlier [25]. These papers where followed by many others, both of physical and mathematical nature [26, 27, 28, 29, 30]. For hyperbolic convective-viscous Cahn-Hilliard equation without source/sink term, i.e. (1.17) with $R\left(X\right)=0$, the exact traveling-wave solution was given in [3].
This paper is organized as follows: in the next Section we give exact traveling wave solution for the Convective-Viscous CHS equation complemented by memory effects. In Section 3 we study the parametric dependence of this solution. In Section 4 we discuss our results.
II Convective Viscous Cahn-Hilliard-Schlögl Equation with memory effects
In the present Section we first give exact traveling-wave solutions for convective viscous Cahn-Hilliard equation [2, 3] supplemented by memory effects and third order reaction terms. So, we first take into account the influence of both external field and dissipation; then we drop the convective and viscous terms setting $\bar{\alpha}=0,\,\,\bar{\eta}=0$, thereby reducing equation to hyperbolic Cahn-Hilliard-Schlögl equation.
To make the calculations somewhat more transparent we presume reaction (1.11) to be irreversible, i.e. $k^{\prime}_{22}=0$ in (1.13). Combining (1.1) and (1.2), substituting the right-hand side of (1.13) into (1.17) for $R\left(X\right)$ we write down the Convective Viscous CHS equation with memory effect, first in terms of the initial variable $X$ (concentration):
@@ II Convective Viscous Cahn-Hilliard-Schlögl Equation with memory effects
Writing down equations (II)-(2.2) we presume implicitly that in the system $A-B-C-X$ the components $A$ and $B$ are in large excess and are not essentially exhausted during the chemical reaction; we have also presumed $M$ to be a constant. Renormalizing $X$, $x^{\prime}$ and $t^{\prime}$, we introduce
$X=uX_{0};\,\,\,x^{\prime}=xL;\,\,\,t^{\prime}=tT.$ (2.3)
Here $X_{0}=\frac{1}{\sqrt{q}}$, $T=\frac{1}{k^{\prime}_{12}X_{0}^{2}}=\frac{q}{k^{\prime}_{12}}$ and $L=\sqrt{MT}=\sqrt{\frac{M}{k^{\prime}_{12}X_{0}^{2}}}=\sqrt{\frac{Mq}{k^{\prime}_{12}}}$. Denoting $\alpha=\bar{\alpha}\frac{X_{0}T}{L}=\bar{\alpha}\frac{1}{\sqrt{k^{\prime}_{12}M}}\,\,$; $\varepsilon^{2}=\frac{\bar{\varepsilon}^{2}}{L^{2}}$; $\eta=\frac{\bar{\eta}}{T}$; $\delta=X_{0}\bar{\delta}=\,\frac{\bar{\delta}}{\sqrt{q}}\,$; $\Theta=\frac{k_{12}A}{k^{\prime}_{12}X_{0}}$, $\Omega=\frac{k_{22}B}{k^{\prime}_{12}X_{0}^{2}}$ and $\tau=\frac{\tau^{\prime}}{T}$, we write down equation (II) in non-dimensional form
@@
Below we presume that the quadratic equation
@@
[…]
$\frac{d}{dz}\left[vu+\alpha\frac{u^{2}}{2}+v\tau u\left(u-u_{1}\right)\left(u-u_{2}\right)\right.$
${\left.+\frac{d}{dz}\left(-v^{2}\tau u-\varepsilon^{2}\frac{d^{2}u}{dz^{2}}+u^{3}-\delta u^{2}-su-v\eta\frac{du}{dz}\right)\right]=u\left(u-u_{1}\right)\left(u-u_{2}\right)}.$ (2.6)
We look for the solution, which connects the stable stationary state of the reaction system $u=u_{2}$ at $z=-\infty$ with the stable stationary state $u=0$ at $z=+\infty$. So, the proper Ansaz for the anti-kink solution (as usually we call ‘‘kinks’’ the solutions with $\frac{du}{dz}>0$, and ‘‘anti-kinks’’ – solutions with $\frac{du}{dz}<0$) with this property will be
$\frac{1}{\kappa}\frac{du}{dz}=u\left(u-u_{2}\right),$ (2.7)
where $\kappa$ is presently unknown positive constant. Then equation (II) could be written as
${\frac{d}{dz}\left[vu+\alpha\frac{u^{2}}{2}\right.-\left(\frac{1}{2\kappa}u^{2}-\frac{u_{1}}{\kappa}u\right)}$
${\left.+\frac{d}{dz}\left(\frac{v\tau}{2\kappa}u^{2}-\frac{u_{1}v\tau}{\kappa}u-\varepsilon^{2}\frac{d^{2}u}{dz^{2}}+u^{3}-\delta u^{2}-\left(v^{2}\tau+s\right)u-v\eta\frac{du}{dz}\right)\right]=0}.$ (2.8)
@@
${+\frac{d}{dz}\left[-\varepsilon^{2}\frac{d^{2}u}{dz^{2}}+u^{3}+\left(\frac{v\tau}{2\kappa}-\delta-v\eta\kappa\right)u^{2}-\left(v^{2}\tau+s+\frac{u_{1}v\tau}{\kappa}-v\eta\kappa u_{2}\right)u\right]=C_{1}},$ (2.9)
@@
$\;\frac{d^{2}u}{dz^{2}}=\kappa^{2}\left(2u^{3}-3u_{2}u^{2}+u_{2}^{2}u\right).$ (2.10)
Regarding the Ansatz (2.7), for the equation (II) to be satisfied the expression under the derivative should be linear in $u$. I.e., for (2.7) to give the solution of (II) two following equations should be satisfied for arbitrary $u$:
@@
where $\beta$ is unknown constant. Setting integration constants equal to zero and equating to zero coefficients at each power of $u$ we finally obtain five constraints on the parameters:
@@
If the constraints (2.13)-(2.17) are fulfilled, the solution of (2.7) is simultaneously solution of (II). Integrating (2.7) and taking $z=0$ for maximal steepness point, we get
$u=\frac{u_{2}}{2}\left[1-\tanh\left(\frac{u_{2}}{2\sqrt{2}\varepsilon}\left(x-vt\right)\right)\right].$ (2.18)
The functional form of solution is rather simple; however, the dependence of the parameters of solution on the system parameters is quite complicated, see [1]. In the next Section we will, following the lines of [1], consider the changes introduced by memory.
III The parametric dependence of solution
Similarly to [1] there are five constraints, and only three unknowns $\kappa,\,v$ and $\beta$. I.e. for the constant velocity transition front to exist, two additional constraints on the values of the stationary states of the reaction system and on the values of the equilibrium states for the phase transition should be imposed. We will presume, as in [1], that the parameters related to reaction system are ‘‘basic’’. Evidently, (2.13)-(2.14) and (2.16), and correspondently $\kappa,\,\,v,\,\,\beta$ coincide with the given in Part I [1] (we remind that notations are different):

### v2

# Cahn–Hilliard model with Schlögl reactions: interplay of equilibrium and non-equilibrium phase transitions. II. Memory effects

@@ Abstract
The present work is a continuation of our previous paper [Condens. Matter Phys., 2020, 23, 33602: 1–17]. It is devoted to the modelling of the interplay of equilibrium and non-equilibrium phase transitions. The modelling of equilibrium phase transition is based on the modified Cahn–Hilliard equation. The non-equilibrium phase transition is modeled by the Second Schlögl reaction system. We consider the advancing front, which combines these both transitions. Different from the first article, we consider here the memory effects, i.e., the effects of non-Fickian diffusion. The traveling wave solution is obtained, and its dependence on the model parameters is studied in detail. The relative importance of memory effects for different process regimes is estimated.
@@
Key words: phase transition, nonequilibrium phase transition, Cahn–Hilliard equation, Schlögl reactions, memory effect, travelling wave
Abstract
Öÿ ðîáîòà є ïðîäîâæåííÿì íàøî¿ ðîáîòè [Condens. Matter Phys., 2020, 23, 33602: 1–17]. Âîíà ïðèñâÿчåíà ìîäåëþâàííþ âçàєìîäi¿ ðiâíîâàæíèõ i íåðiâíîâàæíèõ ôàçîâèõ ïåðåõîäiâ. Ìîäåëþâàííÿ ðiâíîâàæíîãî ôàçîâîãî ïåðåõîäó áàçóєòüñÿ íà ìîäèôiêîâàíîìó ðiâíÿííi Êàíà–Õiëüÿðäà. Íåðiâíîâàæíèé ôàçîâèé ïåðåõiä ìîäåëþєòüñÿ äðóãîþ ñèñòåìîþ õiìiчíèõ ðåàêöié Øëüîãëÿ. Ðîçãëÿäàєòüñÿ íàñòóïàþчèé ôðîíò, ÿêèé ïîєäíóє öi îáèäâà ïåðåõîäè. Íà âiäìiíó âiä ïåðøî¿ ñòàòòi, òóò ìè ðîçãëÿäàєìî åôåêòè ïàì’ÿòi, òîáòî åôåêòè íåôiêêiâñüêî¿ äèôóçi¿. Îòðèìàíî ðîçâ’ÿçîê áiæóчî¿ õâèëi, äåòàëüíî äîñëiäæåíî éîãî çàëåæíiñòü âiä ïàðàìåòðiâ ìîäåëi. Îöiíåíî âiäíîñíó âàæëèâiñòü åôåêòiâ ïàì’ÿòi äëÿ ðiçíèõ ðåæèìiâ ïðîöåñó.
Ключовi слова: ôàçîâèé ïåðåõiä, íåðiâíîâàæíèé ôàçîâèé ïåðåõiä, ðiâíÿííÿ Êàíà–Õiëüÿðäà, ðåàêöi¿ Øëüîãëÿ, åôåêò ïàì’ÿòi, áiæóчà õâèëÿ
1 Introduction
@@
The present work is a continuation of our paper [1]. It was devoted to the modelling of the interplay of equilibrium and non-equilibrium phase transitions. We considered the advancing fronts which “combine”, in some sense, these both transitions. The equilibrium phase transition was modeled on the basis of convective-viscous Cahn–Hilliard equation [2, 3]. The non-equilibrium phase transition was modeled by canonical Schlögl chemical reaction systems [4]. The insight into the history and existing modifications of the Cahn–Hilliard equation was given in [1] and will not be repeated here. We also refer to the pioneering papers and excellent reviews [5, 6, 7, 8]. Below, in the introduction, we first give only a brief explanation of some basic assumptions and a description of the models, introduced in [1]; then we outline the specificities of the present work.
@@
As it was mentioned in [1], the basic underlying idea of Cahn–Hilliard model is that for an inhomogeneous system, e.g., system undergoing a phase transition, the thermodynamic potential (e.g., free energy) should depend not only on the order parameter $X$, but also on its gradient. For an inhomogeneous system, the local chemical potential $\mu$, defined as variational derivative of the thermodynamic potential functional, is as follows:
$\mu=-\bar{\varepsilon}^{2}\Delta X+f(X).$ (1.1)
@@
Herein our notations differ from notations in [1]. $\bar{\varepsilon}$ is usually assumed to be proportional to the capillarity length, and $f\left(X\right)={\mathrm{d}\Phi\left(X\right)}/{\mathrm{d}X}$, where $\Phi\left(X\right)$ is a homogeneous part of the thermodynamic potential. As in [1], in the present communication we take $f\left(X\right)$ in the form of the cubic polynomial (corresponding to the fourth-order polynomial for a homogeneous part of thermodynamic potential):
@@
In this phenomenological model, we always consider the isothermal situation. So we do not show the temperature dependence of the coefficients in (1.2) explicitly. However, if we want to model the approach to a critical state, for such a model the approach to critical temperature will be manifested by merging the stationary states together [1], i.e., by two non-zero roots of the right-hand side of (1.2) approaching the third zero root.
@@
yields, with expression (1.1) for the chemical potential, instead of the usual second order diffusion equation, a fourth-order PDE for the order parameter $X$:
@@
In the presence of the external field, the convective term is added to equation (1.5), and to account for dissipation the viscous term is added to the chemical potential $\mu$ [10, 11, 12, 13, 2, 3]:
@@
Here, $\bar{\alpha}$ is proportional to the field and $\bar{\eta}$ is viscosity; we consider the one-dimensional problem herein.
@@
Among numerous modifications of the Cahn–Hilliard (CH) equation we are interested here in the models, where the nonlinear sink/source terms, e.g., due to a chemical reaction, are inserted into the right-hand sides of (1.4) (and, correspondingly of (1.5) as well):
$\frac{\partial X}{\partial t^{\prime}}=\nabla\left[M\nabla\mu\right]+R\left(X\right).$ (1.7)
For the classic CH equation such a study was pioneered by Huberman [14], Cohen and Murray [15]; see also [16, 17, 18, 19, 20, 21].
@@
General observation is that the presence of chemical reaction can visibly influence the equilibrium phase transition, e.g., freeze the spinodal decomposition or coarsening, stabilizing some stationary inhomogeneous state. On the other hand, the canonical models for non-equilibrium phase transitions in chemical reaction systems were introduced by Schlögl [4]; here, the different “phases” correspond to different stationary states of the system. Schlögl considered two reaction systems, the so-called “First Schlögl Reaction” and the “Second Schlögl Reaction”. The first reaction exhibits a non-equilibrium phase transition of the second order, the second one exhibits a phase transition of the first order (for the details see [4]). If the system simultaneously undergoes an equilibrium phase transition accompanied by a phase separation, it could be of considerable interest to study the interaction of an equilibrium and non-equilibrium phase transitions. We also call such a process “a combined transition” for brevity.
@@
In [1] we considered the convective-viscous Cahn–Hilliard equation (1.6), see [2, 3], complemented by source/sink term $R\left(X\right)$. We considered $R\left(X\right)$ corresponding both to first and second Schlögl reactions. We called these modifications Cahn–Hilliard–Huberman–Cohen–Murray (CHHCM) and Cahn–Hilliard–Schlögl (CHS) equations, respectively. We obtained exact travelling-wave solutions for these modifications. For convective-viscous equation with the first Schlögl reaction, the conditions of simultaneous equilibrium and non-equilibrium phase transitions are very restrictive and rigid. Even more, the equation for the travelling wave becomes degenerate in the absence of an applied field, splitting into second and first order equations. Thus, the constant-velocity combined-transition-front model is not very instructive in this case. On the other hand, for the CHS, the effect of the non-equilibrium transition, i.e., of the reactive system, is much stronger. The transition front may be stopped, or even reversed both by changing the stationary states of the reaction system and by the field. Thus, exploring the “memory effects”, see below, in the present work we consider the convective-viscous CHS only.
The description of the “Second Schlögl Reaction” is given in [1], equations (1.10), (1.11). For the second Schlögl reaction in the absence of diffusion, the evolution of $X$ is described by
$\frac{\mathrm{d}X}{\mathrm{d}t^{\prime}}=\,-k^{\prime}_{12}X^{3}+k_{12}AX^{2}-k_{22}BX+k^{\prime}_{22}C.$ (1.8)
We denote the concentrations with the same letters as species; the concentrations of species $A$, $B$ and $C$ (which are called the “reservoir reagents”) are assumed to be constant and only concentration of $X$ can vary with time and space. We denote the rate constants by $k_{ij}$, $k^{\prime}_{ij}$ for the forward and reverse reactions, respectively; to resemble with notations in [1], we keep the second lower index “2” for the second Schlögl reaction. In [1] we considered convective viscous CH equation with $R(X)$ equal to the right-hand-side of (1.8).
@@
However, in all models described above, the diffusional flux was assumed to be proportional to the gradient of the chemical potential. This “modified” Fick’s law (the original Fick’s law presumes proportionality to the gradient of concentration) was often criticized for the infinite speed of the spread of a diffusing substance. The most popular alternative is the Maxwell-Cattaneo approach, see [22] for a comprehensive discussion; here, we give only the one-dimensional formulae, in accord with the spirit of the present paper. In Maxwell-Cattaneo approach, the mass-conservation, or continuity not changed. However, instead of (1.3) the following relation is proposed:
@@
$J=-\int_{0}^{t^{\prime}}\left(\frac{M}{\tau^{\prime}}\frac{\partial\mu}{\partial x^{\prime}}\right)\exp\left(\frac{t^{\prime\prime}-t^{\prime}}{\tau^{\prime}}\right)\mathrm{d}t^{\prime\prime}.$ (1.10)
Thus, this approach is also called “diffusion with memory effects”. Correspondingly, $\tau^{\prime}$ is considered to be a characteristic time of the memory. On the other hand, eliminating $J$ from (1.9) yields the “hyperbolic modification” of the diffusion equation with a source
@@
Similarly, a hyperbolic modification of the convective-viscous Cahn–Hilliard equation (1.6) with source/sink term takes the form
@@
The hyperbolic modification of the classical Cahn–Hilliard equation was proposed in [23] to model a rapid spinodal decomposition in a binary alloy. However, from purely mathematical point of view — as a singular perturbation of the classic Cahn–Hilliard equation — it was considered even earlier [24]. These papers were followed by many others, both of physical and mathematical nature [25, 26, 27, 28, 29]. For hyperbolic convective-viscous Cahn–Hilliard equation without source/sink term, i.e., (1.12) with $R(X)=0$, the exact travelling-wave solution was given in [3].
This paper is organized as follows: in the next section we give an exact travelling wave solution for the convective-viscous CHS equation complemented by memory effects. In section 3 we study the parametric dependence of this solution. In section 4 we discuss our results.
2 Convective viscous Cahn–Hilliard–Schlögl equation with memory effects
In the present section we first give exact travelling-wave solutions for convective viscous Cahn–Hilliard equation [2, 3] supplemented by memory effects and third order reaction terms. Thus, we first take into account the influence of both external field and dissipation; then we drop the convective and viscous terms setting $\bar{\alpha}=0$, $\bar{\eta}=0$, thereby reducing the equation to hyperbolic Cahn–Hilliard–Schlögl equation.
To make the model and calculations somewhat more transparent we assume the reaction to be irreversible, i.e., $k^{\prime}_{22}=0$ in (1.8). Substituting the right-hand side of (1.8) into (1.12) for $R\left(X\right)$, we write down the convective viscous CHS equation with the memory effect, first in terms of the initial variable $X$ (concentration):
@@ 2 Convective viscous Cahn–Hilliard–Schlögl equation with memory effects
Writing down equations (2)–(2.2) we implicitly assume that in the system $A-B-C-X$ the components $A$ and $B$ are in large excess and are not essentially exhausted during the chemical reaction; we also presumed $M$ to be a constant. Renormalizing $X$, $x^{\prime}$ and $t^{\prime}$, we introduce
$X=uX_{0};\quad x^{\prime}=xL;\quad t^{\prime}=tT.$ (2.3)
Here, $X_{0}={1}/{\sqrt{q}}$, $T={1}/{k^{\prime}_{12}X_{0}^{2}}={q}/{k^{\prime}_{12}}$ and $L=\sqrt{MT}=\sqrt{{M}/{k^{\prime}_{12}X_{0}^{2}}}=\sqrt{{Mq}/{k^{\prime}_{12}}}$. Denoting $\alpha=\bar{\alpha}({X_{0}T}/{L})=\bar{\alpha}({1}/{\sqrt{k^{\prime}_{12}M}})$; $\varepsilon^{2}={\bar{\varepsilon}^{2}}/{L^{2}}$; $\eta={\bar{\eta}}/{T}$; $\delta=X_{0}\bar{\delta}={\bar{\delta}}/{\sqrt{q}}$; $\Theta={k_{12}A}/{k^{\prime}_{12}X_{0}}$, $\Omega={k_{22}B}/{k^{\prime}_{12}X_{0}^{2}}$ and $\tau={\tau^{\prime}}/{T}$, we write down equation (2) in non-dimensional form
@@
Below we assume that the quadratic equation
@@
always has real roots $u_{1}$, $u_{2}$, $0<u_{1}\leqslant u_{2}$, corresponding to stationary states of the reactions system. I.e., $\Theta^{2}-4\Omega\geqslant 0$ which means, in terms of the parameters of the reaction system, $\left(k_{12}A\right)^{2}\geqslant 4k^{\prime}_{12}k_{22}B$. This is simply the necessary condition of the multiplicity of stationary states in Schlögl reaction model, see equations (1.10), (1.11) of [1], and equation (1.8), i.e., the existence of a non-equilibrium phase transition.
Looking for the traveling wave solutions of (2), we introduce the travelling wave coordinate $z=x-vt$. This yields
$\frac{\mathrm{d}}{\mathrm{d}z}\left[vu+\alpha\frac{u^{2}}{2}+v\tau u\left(u-u_{1}\right)\left(u-u_{2}\right)\right.$
${\left.+\frac{\mathrm{d}}{\mathrm{d}z}\left(-v^{2}\tau u-\varepsilon^{2}\frac{\mathrm{d}^{2}u}{\mathrm{d}z^{2}}+u^{3}-\delta u^{2}-su-v\eta\frac{\mathrm{d}u}{\mathrm{d}z}\right)\right]=u\left(u-u_{1}\right)\left(u-u_{2}\right)}.$ (2.6)
We look for the solution, which connects the stable stationary state of the reaction system $u=u_{2}$ at $z=-\infty$ with the stable stationary state $u=0$ at $z=+\infty$. Thus, the proper Ansatz for the anti-kink solution (as usually we call “kinks” the solutions with ${\mathrm{d}u}/{\mathrm{d}z}>0$, and “anti-kinks” — the solutions with ${\mathrm{d}u}/{\mathrm{d}z}<0$) having this property will be
$\frac{1}{\kappa}\frac{\mathrm{d}u}{\mathrm{d}z}=u\left(u-u_{2}\right),$ (2.7)
where $\kappa$ is presently unknown positive constant. Assuming that the solutions of (2.7) are simultaneous solutions of equation (2), we can rewrite (2) as follows:
${\frac{\mathrm{d}}{\mathrm{d}z}\left[vu+\alpha\frac{u^{2}}{2}\right.-\left(\frac{1}{2\kappa}u^{2}-\frac{u_{1}}{\kappa}u\right)}$
${\left.+\frac{\mathrm{d}}{\mathrm{d}z}\left(\frac{v\tau}{2\kappa}u^{2}-\frac{u_{1}v\tau}{\kappa}u-\varepsilon^{2}\frac{\mathrm{d}^{2}u}{\mathrm{d}z^{2}}+u^{3}-\delta u^{2}-\left(v^{2}\tau+s\right)u-v\eta\frac{\mathrm{d}u}{\mathrm{d}z}\right)\right]=0}.$ (2.8)
@@
${+\frac{\mathrm{d}}{\mathrm{d}z}\left[-\varepsilon^{2}\frac{\mathrm{d}^{2}u}{\mathrm{d}z^{2}}+u^{3}+\left(\frac{v\tau}{2\kappa}-\delta-v\eta\kappa\right)u^{2}-\left(v^{2}\tau+s+\frac{u_{1}v\tau}{\kappa}-v\eta\kappa u_{2}\right)u\right]=C_{1}},$ (2.9)
@@
$\frac{\mathrm{d}^{2}u}{\mathrm{d}z^{2}}=\kappa^{2}\left(2u^{3}-3u_{2}u^{2}+u_{2}^{2}u\right).$ (2.10)
Regarding the Ansatz (2.7), for the equation (2) to be satisfied, the expression under the derivative should be linear in $u$. I.e., for (2.7), to get a solution of (2), two following equations should be identically satisfied for arbitrary $u$:
@@
where $\beta$ is an unknown constant. Equating to zero coefficients at each power of $u$, including $C_{1}$, $C_{2}$ as zero power coefficients, we finally obtain five constraints on the parameters:
@@
If the constraints (2.13)–(2.17) are fulfilled, the solution of (2.7) is simultaneously solution of (2). Integrating (2.7) and taking $z=0$ for maximal steepness point, we get
$u=\frac{u_{2}}{2}\left\{1-\tanh\left[\frac{u_{2}}{2\sqrt{2}\varepsilon}\left(x-vt\right)\right]\right\}.$ (2.18)
The functional form of a solution is rather simple; however, the dependence of the parameters of solution on the system parameters is quite complicated, see [1]. In the next section we, following the lines of [1], consider the changes introduced by the memory.
3 The parametric dependence of solution
Similarly to [1] there are five constraints, and only three unknowns $\kappa,\,v$ and $\beta$. I.e., for the constant velocity transition front to exist, two additional constraints on the values of the stationary states of the reaction system and on the values of the equilibrium states for the phase transition should be imposed. We assume, as in [1], that the parameters related to the reaction system are “basic”. Evidently, (2.13)–(2.14) and (2.16), and correspondently $\kappa$, $v$, $\beta$ coincide with the given in Part I [1] (we remind that notations are different):